OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)?

OpenStudy (michele_laino):

Please, you have to apply this formula: \[slope=\frac{ \Delta y }{\Delta x }\]

OpenStudy (anonymous):

\[y-y_1 = m(x-x_1)\] \[(y-y_1)/(x-x_1) = m\] m = slope of line (x,y) and (x1,y1) are the points

OpenStudy (anonymous):

Thanks everyone

OpenStudy (dan815):

-2

OpenStudy (anonymous):

thanks

OpenStudy (dan815):

theyre the ones that did the work